Brian works on commission. For each subscription he sells in a day, Brian earns $20. The function f(x)=20x models this, where x is subscriptions sold and f(x) is the number of dollars Brian earn
A. Is the domain of f(x) all real numbers?
B. List five ordered pairs with x < 10 that satisfy the function?
C. Brian’s boss starts paying him $40 per day in addition to $20 per subscription. How would the graph of f(x) change?
D. Suppose that instead of paying $40 extra per day, Brian’s boss decides to pay him $30 per subscription. How would the graph of f(x) change?
1
Expert's answer
2019-09-11T13:32:05-0400

(a) "f(x)=20x \ \"

Here "x" represent the number of subscriptions sold which are natural numbers.

Hence as per the question ,domain of "f(x)" is "x \in N"

(b) Five ordered pairs for "x<10"

"1. x=1;f(x)=20 x=20 \times 1=20 \implies(x,f(x))=(1,20)"

"2. x=2;f(x)=20 x=20 \times 2=40 \implies(x,f(x))=(2,40)"

"3. x=3;f(x)=20 x=20 \times 3=60 \implies(x,f(x))=(3,60)"

"4. x=4;f(x)=20 x=20 \times 4=80 \implies(x,f(x))=(4,80)"

"5. x=5;f(x)=20 x=20 \times 5=100 \implies(x,f(x))=(5,100)"

(c) Initial model"=f(x)=20 x"

It is a straight line passing through origin with slope +20.

When boss starts paying $40 extra,

the new model will be

"f(x)=20x+40"

The new graph will also be a straight line with slope 20 but it does not passes through origin.

It has a "y-axis" intercept of +20 which means he will get $40 even if he does not sell any subscriptions in a day.

(d) New model will be "f(x)=30 x"

It is a straight line passing through origin with slope +30 bisecting the first and third quadrant.